OpenStudy (anonymous):

if alpha,beta are the zeroes of the polynomial p(x)=x^2+x-6,then the value of 1/alpha^2+1/beta^2.

terenzreignz (terenzreignz):

\[x^{2} + x - 6 = (x + 3)(x - 2)\]

OpenStudy (anonymous):

cn u plzzzz solve i wanna c i didn't gt d answer

OpenStudy (anonymous):

-3 and 2 are the zeroes

terenzreignz (terenzreignz):

Equate it to 0 (x + 3)(x - 2) = 0 Then either x = -3 or x = 2 2 and -3 are your alpha and beta :)

OpenStudy (anonymous):

so 1/\[\alpha^2\] +1/\[\beta^2\]=????????

OpenStudy (anonymous):

1/9+1/4

OpenStudy (anonymous):

gt d answer thx

OpenStudy (shubhamsrg):

well yes you can do it like this : there is another method : 1/a^2 + 1/b^2 = (a^2 + b^2)/(ab)^2) = ((a+b)^2 - 2ab)/(ab)^2) here a =alpha,,b = beta.. just use the sum of roots and product of roots value,,you should get the same ans.